OpenStudy (anonymous):

Braden improves his score with each practice quiz he takes. The number of incorrect answers can be represented by the expression 20 – 5n where n represents the number of practice quizzes he has taken. How many answers will he get incorrect after taking 2 practice quizzes? 15 13 10 5

OpenStudy (anonymous):

Can you substitute n for two in the given formula and solve?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

You have the formula 20 - 5n, where n is the number of tests taken. The number of tests taken are 2. So replace n in the formula with 2, and solve it. That it is say, change it from 5*n to 5*2.

OpenStudy (anonymous):

and 5 times 2 is 10 so the answer is 10?

OpenStudy (anonymous):

Yes! 20 - 5n = 20 - 5*2 = 20 - 10 = 10.

OpenStudy (anonymous):

tnx so much

OpenStudy (anonymous):

No worries.